OpenStudy (anonymous):

lim as x approaches 1 of (x^2-1/ x-1 )

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} \frac{x^2-1}{x-1}= \lim_{x \rightarrow 1} \frac{(x-1)(x+1)}{x-1}= (1+1)=2\]

OpenStudy (anonymous):

hm i do it in same way dear